# Evaluating Human Alignment and Model Faithfulness of LLM Rationale

### v1

@@ 4.1 Human Alignment
The annotated rationales provide explanations for the ground truth label. Therefore, for the evaluation of human alignment, we first request the model to provide a rationale for the provided label. With this, we create an array of binary values where for each word in the input sentence, we indicate a 1 if it is present among the generated words (0 otherwise). By comparing these vectors with the equivalent binary representations of human annotations, we calculate the F1 score. This F1 score for human alignment is reported in Table 2 for the e-SNLI and MedicalBios datasets.
Figure 2: Accuracy changes throughout 5 epochs of fine-tuning. (PT denotes the pre-trained model’s accuracy)
Firstly, prompting-based methods outperform attribution-based methods on human alignment. Short or normal prompting demonstrates superior performance compared to attribution-based methods in nearly all datasets and models except for Llama-3-8B on e-SNLI, which is also comparable. This gap can be attributed to the reliance of attribution-based methods on the classification capability of LLMs (which might be subpar).
@@
Secondly, we note that providing additional information about the number of words selected by humans in Top-Var settings enhances alignment, indicating disparities between model thresholds for word importance in Unbound prompting compared to human annotators. Furthermore, the random baseline, which involves selecting Top-Var random words in each sentence, yields F1 scores of 0.27±0.04 and 0.22±0.01 for e-SNLI and MedicalBios respectively across 100 seeds. Contrasted with the Top-Var rows in Table 2, this indicates that both attribution-based and prompting-based approaches exhibit superior alignment compared to a random baseline.
@@
Figure 3: Human alignment F1$\uparrow$ changes throughout 5 epochs of fine-tuning. (PT denotes the pre-trained model)
Thirdly, across models, the performance comparison between normal and short prompts is varying and inconclusive. For example, short prompts perform better for Mistral, LLama-3, and GPT-4-Turbo, while they perform worse for LLama-2 and GPT-3.5 models.
Finally, comparing the evaluated models reveals GPT-4-Turbo to be the most aligned with humans. However, other models demonstrate task-dependent alignment, with some excelling in e-SNLI and others in MedicalBios.
Model Mistral-7B FT Instruct-v0.2 Llama-2-7b FT chat Llama-3-8B FT Instruct
Dataset E-SNLI MedicalBios E-SNLI MedicalBios E-SNLI MedicalBios
Method Selection
Attribution-based
Saliency Top-Ratio 37.36 (+0.20) 27.29 (−1.69) 36.34 (+7.71) 31.73 (+0.38) 34.76 (+2.57) 30.18 (+2.38)
Input×Gradient Top-Ratio 33.71 (+2.30) 27.47 (−0.04) 36.98 (+10.13) 32.64 (+1.68) 38.59 (+2.27) 31.91 (+2.08)
Saliency Top-Var 44.87 (−1.19) 37.12 (−1.18) 45.63 (+8.37) 42.80 (+1.31) 42.90 (+3.13) 39.57 (+3.21)
Input×Gradient Top-Var 40.95 (+2.09) 35.67 (−1.41) 46.01 (+9.36) 41.91 (+1.46) 46.81 (+2.06) 43.31 (+3.83)
Prompting-based
Normal Prompt Unbound 40.48 (+1.53) 41.66 (−0.84) 45.28 (−0.17) 47.26 (−1.89) 39.04 (−0.54) 48.41 (−0.64)
Short Prompt Unbound 39.49 (−2.35) 42.74 (+0.48) 43.11 (−0.43) 43.45 (−0.29) 43.86 (+0.26) 50.45 (−0.83)
Normal Prompt Top-Ratio 40.99 (−0.17) 41.59 (+0.43) 44.40 (−1.50) 43.29 (−0.70) 35.42 (−0.56) 47.00 (−0.58)
Short Prompt Top-Ratio 42.01 (+0.18) 43.91 (+1.85) 43.23 (+2.33) 45.29 (−0.36) 38.15 (+0.69) 47.68 (+0.77)
Normal Prompt Top-Var 45.70 (−0.24) 48.51 (+0.08) 50.58 (−0.35) 51.35 (−0.88) 41.69 (+0.02) 59.70 (+0.85)
Short Prompt Top-Var 48.35 (+1.38) 51.96 (+3.88) 49.43 (+2.09) 52.43 (+1.89) 45.57 (+0.92) 58.57 (+0.41)
Table 3: Human alignment F1$\uparrow$ score in E-SNLI and MedicalBios datasets after fine-tuning for 5 epochs on the respective dataset. The difference in alignment between the fine-tuned and pre-trained model (Table 2) is reported in the parentheses. (Changes greater than 1.5% are denoted using a bold font.)
4.2 Effect of Fine-tuning on Alignment
As described by Wang et al. (2024) and Zhong et al. (2023), zero-shot LLMs may underperform small fine-tuned models such as BERT. And smaller LLMs like LLaMA-2-7B might even collapse entirely. We observe similar failure patterns, where models generate a single label from the possible options regardless of the input sentence. This phenomenon is illustrated in Figure 2, where the pre-trained (PT) open LLMs achieve near-random accuracy (33%) on the e-SNLI dataset. This issue raises the question of whether fine-tuning these models to improve their classification performance also aids in aligning their explanations more closely with human expectations.
To address this issue, we fine-tune the LLMs using LoRA Hu et al. (2022), a parameter-efficient fine-tuning technique. The hyperparameters for fine-tuning are provided in Table 7. Figure 2 demonstrates that the classification performance of LLaMA-2, LLaMA-3, and Mistral improves significantly after fine-tuning on the e-SNLI dataset and shows slight improvements on the MedicalBios dataset, outperforming GPT-4-Turbo in both cases.
We rerun the human alignment experiments for “Short Prompt Top-Var” and “Input×Gradient Top-Var” across all epochs, as shown in Figure 3. The results suggest a general trend of improved alignment of both methods with increasing epochs.
To analyze the generalization of this trend more comprehensively, we conduct alignment experiments on all prompting and attribution methods in Table 3 on the final epoch (5th epoch) of fine-tuning. The difference in human alignment between the fine-tuned and pre-trained models (Table 2) is reported in parentheses.
@@ 4.2 Effect of Fine-tuning on Alignment
Overall, the results show more positive change than negative in alignment. Among the prompting-based methods, “Short Prompt Top-Var” demonstrates the highest gains from fine-tuning. Attribution-based methods, particularly in LLaMA models, exhibit significant alignment improvements after fine-tuning. Moreover, fine-tuning can guide the model’s attention to the correct words, especially in datasets like e-SNLI where pre-trained classification accuracy was low. We have demonstrated qualitative examples of such cases in Figure 5. As a result, these gradient-based methods can identify more human-aligned rationales by tracing back the attributions from the output label to the input sentence in fine-tuned models.
Model Mistral-7B Instruct-v0.2 Llama-2-7b chat Llama-3-8B Instruct GPT-3.5 Turbo 1106
Dataset E-SNLI MedBios E-SNLI MedBios E-SNLI MedBios E-SNLI MedBios
Masking Part Method Selection
Attribution-based
Input Saliency Top-Ratio 3.00 50.00 2.36 31.82 29.33 36.11 - -
Input×Gradient Top-Ratio 2.33 44.32 2.02 29.09 26.67 40.74 - -
Prompting-based
Input Normal Prompt Unbound 3.33 59.09 1.01 30.91 38.67 41.67 77.00 41.82
Short Prompt Unbound 3.67 59.09 1.68 35.45 44.00 47.22 73.67 37.27
Normal Prompt Top-Ratio 1.67 28.41 1.68 33.64 24.00 45.37 41.33 24.55
Short Prompt Top-Ratio 2.33 30.68 1.01 26.36 22.67 46.30 44.33 20.91
Baselines
Input Human Human 4.00 60.23 2.02 38.18 31.67 43.12 55.67 36.36
Random Random 1.33 9.09 1.01 10.91 24.67 14.68 41.33 5.45
Everything Everything 2.33 97.73 0.34 60.00 69.67 70.64 89.00 74.55
@@ 4.3 Faithfulness to the Model
While human alignment provides a useful measure of the plausibility of LLM rationales, it is also important to consider the faithfulness of these rationales to the model’s actual decision-making process. A word may be crucial for the model’s decision even if it does not align with human rationale and vice versa. Therefore, we must ask: Are the self-explanations genuinely influential in the model’s decision-making process?
@@ 4.3.1 Limitations of Faithfulness Evaluation before Fine-Tuning
Table 4 presents the faithfulness flip rate of the pre-trained LLMs. A noteworthy finding is that in the e-SNLI dataset, where classification accuracy was notably low (Figure 2), both attribution-based and prompting-based methods resulted in a very small flip rate. Even more concerning, masking all the words in the input sentence led to less than a 3% flip rate for the Mistral and LLaMA-2 models (Mask Everything).
In further exploration, Figure 4 illustrates the Input×Gradient attributions of the predicted label, shown in green, to all instruction and input words, shown in shades of red. We notice that in the pre-trained Llama-2 model, the prediction for the label “entailment” is incorrect, with the model placing excessive emphasis on the word “entailment” in the instruction while largely ignoring the input sentence including the premise and hypothesis sentences. However, after fine-tuning, the attribution distribution becomes less skewed, leading to a correct prediction by the model.
Model Mistral-7B FT Instruct-v0.2 Llama-2-7b FT chat Llama-3-8B FT Instruct
Dataset E-SNLI MedicalBios E-SNLI MedicalBios E-SNLI MedicalBios
Selection Method
Unbound Normal Prompt 64.00 34.51 67.00 17.70 64.33 18.75
Short Prompt 53.00 37.17 67.33 21.24 72.00 25.00
Top-Ratio Attention 50.33 23.89 69.67 14.16 49.33 5.36
Saliency 51.00 23.89 59.67 21.24 60.00 17.86
Input×Gradient 45.33 21.24 61.33 22.12 58.00 18.75
Normal Prompt 35.33 17.70 52.33 19.47 45.00 20.54
Short Prompt 40.33 17.70 60.33 23.89 51.33 20.54
Top-Var Attention 48.00 20.35 61.00 11.50 48.67 5.36
Saliency 52.67 14.16 54.67 17.70 55.00 14.29
Input×Gradient 45.33 19.47 53.33 19.47 56.00 15.18
Normal Prompt 37.00 15.04 50.33 16.81 47.00 16.07
Short Prompt 40.33 19.47 55.00 17.70 50.67 16.07
Baselines
Top-Var Human 49.67 21.24 60.00 24.78 62.00 24.11
Top-Var Random 32.67 9.73 45.00 7.96 36.67 6.25
Everything Everything 66.33 84.07 63.00 75.22 77.33 74.11
[…]
Therefore, we hypothesize that the pre-trained model focuses more on the instruction rather than the input sentence. Moreover, in Table 10, we investigate two masking scenarios. The first scenario, denoted as Input, involves masking only the words from the input sentence (e.g., the premise and hypothesis in e-SNLI) while leaving the instruction intact, similar to Table 4. The second scenario, denoted as Input&Instruction, extends masking to the entire instruction and input, constituting the entire prompt. When we extend the masking to include the instruction, the flip rate can increase up to 100%. This indicates that the model relies heavily on the instruction for its decisions, regardless of the input sentence. This phenomenon aligns with findings by Yin et al. (2023) and Kung and Peng (2023), who both found that, among all segments of a prompt, label information or output space is essential for the model’s performance. This raises concerns about the reliability of this experiment for measuring model faithfulness in LLMs.
Consequently, we argue that to conduct a more robust faithfulness experiment on LLMs, it is not advisable to solely rely on pre-trained models, as their classification accuracy can vary depending on the model and dataset Madsen et al. (2024). Instead, we suggest aligning the experiment more closely with the scenario of fine-tuned encoder-based models Ferrando et al. (2022); Modarressi et al. (2023) by training the LLMs and assessing faithfulness on the fine-tuned model.
@@ 4.3.2 Faithfulness after Fine-Tuning
Table 5 displays the faithfulness flip rate of the fine-tuned open models on e-SNLI and MedicalBios. The number of masked words can directly influence the prediction flip rate in this experiment. To ensure a fair comparison, we limit the number of words considered in the Top-Ratio and Top-Var selections to the top $k$ words for each sentence across different methods. This approach is particularly important for prompting-based methods, as LLMs struggle to follow instructions involving fine-grained hard constraints Sun et al. (2023).
First, we see that fine-tuning has effectively addressed the near-zero flip rate (Table 4) in e-SNLI, indicating that the model is no longer completely disregarding the input sentence.
Second, a comparison of results in each selection group of “Top-Ratio” and “Top-Var” (with a similar number of masked words) reveals that attribution-based methods generally outperform prompting. Consistent results can be seen in our top-$k$ experiments in the appendix A.1 and Table 8. This difference can be attributed to the fact that attribution methods base their explanations on the model’s internal processes, whereas prompting may provide plausible answers without direct access to this information, potentially diverging from the truth of the model’s inner workings. Additionally, prompting is affected by the model’s ability to follow instructions, which may result in the generation of an inaccurate number of words or the inclusion of words not present in the input sentence, leading to less faithful results.
Third, we also present the flip rate after masking human rationales in Table 5. These rates are comparable to the “Top-Var” selection of other methods, as they involve the same number of masked words. Despite expectations that the model would better recognize the importance of words for its own decisions, prompting methods consistently underperformed human rationales, and attribution methods did so in half of the cases. This result emphasizes that while the current methods demonstrate a degree of faithfulness, there remains room for further refinement and enhancement.
@@ 5 Conclusions
In this study, we investigated the extraction of rationales in Large Language Models (LLMs) with a focus on human alignment and model faithfulness. Our experiments encompassed both prompting-based and attribution-based methods across various LLM architectures and datasets. Before fine-tuning, we observed that prompting generally yielded better human alignment, even when classification performance was poor. However, the reliability of faithfulness evaluations was compromised by low classification performance and collapsing predictions in pre-trained models highlighting the need for refining faithfulness evaluation setup.
To address this, we fine-tuned the models to enhance their accuracy on classification tasks, which led to improvements in aligning their explanations more closely with human expectations. In this scenario, although prompting showed superior alignment before, its faithfulness in reflecting model decision-making was not as strong as that of attribution-based methods.
Despite these improvements, a gap remained between the models’ rationales and human rationales in both alignment and faithfulness. This highlights the need for the development of more advanced explanation methods to bridge this gap.

### v2

@@ 4.2 Human Alignment
The human-annotated rationales provide explanations for the ground truth label. Therefore, we first ask the model to generate its rationale for the true label (by prompting or attribution). Then to measure alignment, we compute the F1 score, which balances precision (the proportion of correctly generated words out of all generated words) and recall (the proportion of correctly generated words out of all relevant words). This score compares the model’s generated rationales with human-annotated ones, assessing how well the model aligns with human explanations. Table 2 presents the F1 scores for the pre-trained and fine-tuned (epoch 10) models.
First, comparing the evaluated models reveals GPT-4-Turbo to be the most aligned with humans. Although the other models show varying levels of alignment depending on the task, GPT-3.5-Turbo and Llama3 often provide better prompting explanations than Llama2 and Mistral.
@@
Second, we note that providing additional information about the number of words selected by humans in Top-Var settings enhances alignment, indicating disparities between model thresholds for word importance in Unbound prompting compared to human annotators.
Third, attribution-based methods tend to align more closely with human reasoning than prompting-based methods, especially after fine-tuning. This suggests that the model may be attending to words in a manner similar to human thought processes, although this cannot be fully explained through language modeling. Furthermore, among the attribution-based methods, Input×Gradient and Saliency appear to outperform raw attention weights, which is anticipated based on existing literature Ferrando et al. (2022).
@@
Figure 3: Accuracy and Human Alignment F1 score of 10 epochs of fine-tuning.
Method Input×Gradient Prompting
Dataset Model
E-SNLI Mistral-7B 0.85 (0.00) -0.35 (0.30)
Llama-2-7b 0.98 (0.00) -0.52 (0.10)
Llama-3-8B 0.88 (0.00) -0.90 (0.00)
FEVER Mistral-7B 0.97 (0.00) -0.06 (0.86)
Llama-2-7b 0.90 (0.00) -0.36 (0.28)
Llama-3-8B 0.70 (0.02) 0.86 (0.00)
MedicalBios Mistral-7B 0.58 (0.06) 0.16 (0.64)
Llama-2-7b -0.09 (0.79) 0.57 (0.07)
Llama-3-8B 0.61 (0.04) 0.61 (0.05)
Table 3: The Pearson Correlation (p-value) of Human alignment F1$\uparrow$ score and classification accuracy across 10 epochs of fine-tuning each model. Input×Gradient is highly correlated with accuracy while prompting doesn’t show a clear trend.
4.3 Effect of Classification Performance on Alignment
To analyze the effect of task performance on alignment more comprehensively, we rerun the alignment experiments for “Prompt Top-Var” and “Input×Gradient Top-Var” across all 10 epochs, as shown in Figure 3. Moreover, Table 3 shows the correlation between the Alignment F1 and the model’s classification accuracy.
@@ 4.3 Effect of Classification Performance on Alignment
The results suggest a general trend of improved alignment for attribution-based methods with increasing accuracies throughout epochs. [^6] With the exception of MedicalBios, which does not show a significant range of accuracy improvement (Less than 10%), making it not a great candidate for this analysis. This enhancement can stem from their reliance on the classification prompt and the classification capabilities of LLMs, which may fall short in off-the-shelf models. Consequently, fine-tuning enhances performance and affects the model’s internal processes, potentially making them more accessible and transparent for detection via attention or gradient-based methods. Moreover, fine-tuning can guide the model’s attention to the correct words, especially in datasets like e-SNLI where pre-trained classification accuracy was low. [^7] Figure A.3 illustrates qualitative examples of such cases. As a result, these gradient-based methods can identify more human-aligned rationales by tracing back the attributions from the output label to the input sentence in fine-tuned models. Nonetheless, in prompting methods, this improvement in classification seems to function independently of the explanation task, as it does not demonstrate clear or significant enhancements.
Pre-trained Faithfulness Fine-tuned Faithfulness
E-SNLI FEVER MedBios E-SNLI FEVER MedicalBios
Model Method
Mistral-7B Instruct-v0.2 Prompting 1.00 1.70 19.27 16.33 (+15.33) 7.33 (+5.63) 8.85 (−10.42)
Attention 0.00 15.65 24.77 43.67 (+43.67) 11.33 (−4.31) 12.39 (−12.38)
Saliency 0.00 14.97 30.28 50.00 (+50.00) 12.33 (−2.63) 15.93 (−14.35)
Input×Grad 0.00 14.29 28.44 41.33 (+41.33) 11.67 (−2.62) 16.81 (−11.63)
Random 0.33 12.24 5.50 27.00 (+26.67) 5.00 (−7.24) 2.65 (−2.85)
Human 0.67 1.70 44.04 50.00 (+49.33) 18.67 (+16.97) 14.16 (−29.88)
All 0.00 1.70 100.00 68.67 (+68.67) 23.00 (+21.30) 84.96 (−15.04)
Llama-2-7b chat Prompting 1.34 0.00 20.56 33.00 (+31.66) 8.33 (+8.33) 14.16 (−6.40)
Attention 0.67 4.00 25.23 31.33 (+30.66) 9.00 (+5.00) 14.16 (−11.07)
Saliency 1.68 4.33 37.38 46.67 (+44.99) 12.33 (+8.00) 16.81 (−20.57)
Input×Grad 2.01 4.33 37.38 46.00 (+43.99) 11.33 (+7.00) 16.81 (−20.57)
Random 1.34 3.33 15.89 26.67 (+25.32) 7.00 (+3.67) 5.31 (−10.58)
Human 1.01 0.00 50.47 50.00 (+48.99) 14.33 (+14.33) 20.35 (−30.11)
All 0.00 0.00 71.03 57.33 (+57.33) 19.33 (+19.33) 72.57 (+1.54)
Llama-3-8B Instruct Prompting 15.33 1.33 36.94 22.33 (+7.00) 11.33 (+10.00) 16.07 (−20.87)
Attention 21.33 17.33 21.62 29.67 (+8.33) 11.33 (−6.00) 10.71 (−10.91)
Saliency 22.00 17.67 27.03 45.00 (+23.00) 12.33 (−5.33) 15.18 (−11.85)
Input×Grad 22.67 16.33 28.83 46.00 (+23.33) 14.00 (−2.33) 18.75 (−10.08)
Random 18.00 12.67 4.50 28.33 (+10.33) 8.67 (−4.00) 0.89 (−3.61)
Human 24.00 1.33 40.54 51.00 (+27.00) 18.00 (+16.67) 22.32 (−18.22)
All 79.00 1.33 67.57 81.33 (+2.33) 29.67 (+28.33) 73.21 (+5.65)
@@ 4.4 Faithfulness to the Model
While human alignment provides a useful measure of the plausibility of LLM rationales, it is more important to consider the faithfulness of these rationales to the model’s actual decision-making process. A word may be crucial for the model’s decision even if it does not align with human rationale and vice versa. Therefore, we must ask: Are the self-explanations genuinely influential in the model’s decision-making process?
@@ 4.4.1 Limitations of Faithfulness Evaluation before Fine-Tuning
Table 4 presents the faithfulness flip rate of the pre-trained and fine-tuned LLMs. A noteworthy finding is that in the e-SNLI and FEVER datasets, where classification accuracy was notably low (Figure 2), both attribution-based and prompting-based methods result in a very small (Less than 5%) flip rate in pre-trained models. Even more concerning, masking all the words in the input sentence led to less than a 2% flip rate for the Mistral and Llama-2 models (Mask All).
This issue arises from pre-trained models being heavily biased toward specific labels in poorly performing datasets (Figure A.1). For instance, in the e-SNLI dataset, Llama2, Llama3, and GPT-3.5 predominantly predict “entailment” in over 80% of the examples. Consequently, even masking the entire input does not change their biased predictions, resulting in extremely low faithfulness flip rates.
Figure 4 illustrates the relationship between accuracy and faithfulness flip rate, based on 10 epochs of fine-tuning each model on each dataset. Additionally, Table 5 displays the correlation between model faithfulness and classification accuracy over the same 10 epochs, alongside the pre-trained model’s accuracy. Both analyses indicate that models and datasets with lower pre-trained performance, relative to the zero-rate baseline, tend to show a stronger correlation between improvements in accuracy and increases in faithfulness. This phenomenon can also be attributed to label bias. When a pre-trained model has very low accuracy, it often indicates significant label bias, meaning that masking any number of words may not impact the biased decision. Fine-tuning the model helps mitigate this label bias (Figure A.1), leading to a higher faithfulness flip rate and, consequently, a correlation with accuracy improvements.
Figure 4: Accuracy and Faithfulness Flip Rate of 10 epochs of fine-tuning.
Although fine-tuning is not a definitive solution, it brings attention to a critical issue in previous faithfulness evaluations of LLMs, which were primarily developed with encoder-based models like BERT in mind. Encoder-based models are typically fine-tuned for specific classification tasks, whereas LLMs are pre-trained for broader tasks such as language modeling and instruction following. As a result, using the same faithfulness evaluations assesses the explanations over language modeling capabilities of LLMs rather than classification tasks, where word attributions can differ significantly. The assumption underlying these evaluations is that if a rationale is truly faithful, masking it should meaningfully change the model’s decision. However, our research shows that for LLMs such as Llama2 and datasets like e-SNLI, the model’s performance can be so poor (e.g., consistently predicting the same class) that even masking the entire input text has little to no impact on its predictions, leading to extremely low faithfulness. This problem is further exacerbated by the differences between BERT’s straightforward classifier approach and the prompting used for LLMs, which introduces additional prompt engineering variations and complicates the evaluation process. This performance issue likely led to conclusions in other papers that faithfulness is highly model- and task-dependent. Thus, our study argues that the traditional BERT-based faithfulness evaluation methods are inadequate for current LLMs. Fine-tuning is one approach to addressing this evaluation issue, nonetheless, further research is needed to develop more accurate methods for evaluating the faithfulness of LLM rationales.
PT Acc. Correlation
Dataset Model
E-SNLI Mistral-7B 31.00 0.68 (0.00)
Llama-2-7b 37.67 0.53 (0.01)
Llama-3-8B 38.00 0.58 (0.00)
FEVER Mistral-7B 62.00 0.19 (0.41)
Llama-2-7b 45.33 0.55 (0.01)
Llama-3-8B 70.33 0.07 (0.77)
MedicalBios Mistral-7B 90.27 -0.60 (0.00)
Llama-2-7b 86.73 -0.65 (0.00)
Llama-3-8B 83.19 -0.91 (0.00)
Table 5: The “Pearson Correlation (p-value)” between Faithfulness Flip Rate and Classification Accuracy over 10 fine-tuning epochs. Pre-trained accuracy ("PT Acc.") is also reported. Models and datasets with lower pre-trained performance show a stronger correlation between faithfulness and accuracy gains.
@@ 4.4.2 Faithfulness after Fine-Tuning
Table 4 also displays the faithfulness flip rate of the fine-tuned models. A comparison of results after fine-tuning (where the near-zero flip rate issue is addressed) reveals that attribution methods outperform prompting methods in faithfulness. This difference can stem from attribution methods basing their explanations on the model’s internal processes, whereas prompting may provide plausible answers without direct access to this information, potentially diverging from the truth of the model’s inner workings. Additionally, prompting is affected by the model’s ability to follow instructions, which may result in the generation of an inaccurate number of words or the inclusion of words not present in the input sentence, leading to less faithful results.
Another finding from Table 4 and Figure 4 is that for MedicalBios, where the model was already performing well, further fine-tuning leads to a decrease in faithfulness. This can occur because additional fine-tuning allows the model to learn more cues from the training set, making it less sensitive to word masking. As a result, masking the same number of words may no longer flip its predictions, since the model can rely more effectively on signals from the unmasked portions.
Finally, we present the flip rate after masking human rationales in Table 4. Despite expectations that the model would better recognize the importance of words for its own decisions, these methods consistently underperform human rationales. This result emphasizes that while the current methods demonstrate a degree of faithfulness, there remains room for further refinement and enhancement.
@@ 5 Conclusions
In this study, we investigated extracting rationales from LLMs, focusing on human alignment and model faithfulness. We experimented with prompting and attribution methods across different LLM architectures and datasets. Our study shows that attribution methods generally provide rationales that align more closely with human reasoning compared to prompting methods, especially after fine-tuning. Of greater importance, attribution methods also outperform prompting approaches in faithfulness, as they more accurately reflect the model’s internal decision-making process. Moreover, we showed that fine-tuning reduces label bias and correlates with faithfulness in low-accuracy datasets, explaining the model and task-dependent results of previous works. Despite these improvements, a gap persists between the models’ rationales and human rationales in alignment and faithfulness, underscoring the need for the development of more advanced explanation methods to bridge this gap.